### v1

@@ Abstract
Melanoma is a malignant tumor originating from skin cell lesions. Accurate and efficient segmentation of skin lesions is essential for quantitative medical analysis but remains challenging. To address this, we propose ScaleFusionNet, a segmentation model that integrates Cross-Attention Transformer Module (CATM) and AdaptiveFusionBlock to enhance feature extraction and fusion. The model employs a hybrid architecture encoder that effectively captures both local and global features. We introduce CATM, which utilizes Swin Transformer Blocks and Cross Attention Fusion (CAF) to adaptively refine encoder-decoder feature fusion, reducing semantic gaps and improving segmentation accuracy. Additionally, the AdaptiveFusionBlock is improved by integrating adaptive multi-scale fusion, where Swin Transformer-based attention complements deformable convolution-based multi-scale feature extraction. This enhancement refines lesion boundaries and preserves fine-grained details. ScaleFusionNet achieves Dice scores of 92.94% and 91.65% on ISIC-2016 and ISIC-2018 datasets, respectively, demonstrating its effectiveness in skin lesion analysis. Our code implementation is publicly available at GitHub.
@@ 1 Introduction
The incidence of melanoma has risen significantly in recent decades, driven by increasing environmental pollution and ultraviolet radiation [1]. This trend has drawn significant attention from the global medical community. Early detection is essential for improving treatment outcomes and patient survival. Traditional diagnostic methods, such as clinical observation and tissue biopsy are limited by subjectivity and invasiveness which make them unsuitable for large-scale screening. Medical image segmentation offers a noninvasive and high-precision alternative that gives clinicians with more detailed and accurate information. This technology has the potential to greatly enhance early detection and treatment of skin cancer.
@@
With the rapid advancement of deep learning, convolutional neural network (CNN)-based medical image segmentation models have gained prominence, achieving notable success in tasks such as skin lesion segmentation. Among these, the U-Net model, introduced by Ronneberger et al. [2], stands out as a pioneering framework. U-Net has demonstrated remarkable performance in medical image segmentation and has established a distinctive U-shaped architectural paradigm. However, due to the inherent limitations of convolutional operations, CNN-based methods often struggle to model global contextual information effectively. This limitation becomes particularly evident in medical image segmentation tasks with significant inter-sample variability, such as skin lesion segmentation. To address this challenge, researchers have explored various strategies, including the use of large kernel convolutions, dilated convolutions, and other techniques aimed at expanding receptive fields [3, 4, 5, 6, 7]. For instance, Hu et al. [8] employed self attention to enhance receptive fields, while Tang et al. [9] proposed model, leveraging large convolutional kernels and skip fusion to achieve promising results in tasks like breast nodule ultrasound image segmentation. Inspired by the ConvNeXt model [10], Han et al. [11] developed a medical image segmentation method based on large kernel convolutions, successfully applying it to tasks such as retinal vessel segmentation. Despite these advancements, simply increasing the size of convolutional kernels may not fully resolve the challenge of modeling global features, as the fundamental constraints of the receptive field remain.
@@
Recently, Transformer [12], serving as a feature extraction technique with potent global modeling capabilities, has attained notable success in both natural language processing and computer vision domains. Cai et al. [13] unveiled the BiADATU‐Net, amalgamating Transformer and feature adaptation modules, which resulted in promising outcomes across several publicly accessible skin lesion segmentation datasets. Zhang et al. [14] devised DAE-Former, a pure Transformer U-shaped medical image segmentation model, harnessing efficient Transformers steered by dual attention, akin to Swin-Unet [15], and showcased commendable performance across diverse image segmentation datasets, including ISIC-2018 [16]. Nonetheless, due to the fact that Vision Transformers can only output single-scale feature representations, they lack the ability to capture multi-scale information in two-dimensional images [17, 18]. As a result, Transformer-based medical image segmentation models may struggle to seamlessly integrate multi-scale information, leading to insufficient attention to lesion regions and incomplete decoding of feature details.
@@ 5.1 Results on ISIC-2016
To better compare with current mainstream models, 11 prominent models were selected for comparison with the proposed ScaleFusionNet. Table 1 shows that ScaleFusionNet achieved a DSC score of 92.94% and an IOU score of 87.35% on the ISIC-2016 dataset, which are the average results of five repeated experiments, demonstrating outstanding performance. Compared to the Swin-Unet model, ScaleFusionNet improved by 2.82% in the DSC metric and 4.14% in the IOU metric. When compared to MISSFormer, ScaleFusionNet exhibited enhancements of 2.48% and 3.43% in the DSC and IOU metrics, respectively. Against the D-LKA model, ScaleFusionNet still showed improvements of 0.11% and 0.20% in the DSC and IOU metrics, respectively. This indicates that ScaleFusionNet more accurate than D-LKA to detect the refined boundary of skin lesions. Compared to methods like U-Net, although ScaleFusionNet employs a more complex architecture to address potential issues in U-shaped medical image segmentation models, we find this approach justified given the 5.13% performance improvement and the reduced computational load. In clinical applications, a faster and lighter model can support a broader range of compatible use cases, which is crucial for hospitals or organizations with limited computational resources. Concurrently, to enhance the assessment of model performance, we selects 7 high-performing models for qualitative scrutiny of experimental outcomes, elucidating the differences among them. The red areas in the illustrations represent ground truth labels meticulously annotated by experts, reflecting clinical doctors’ diagnostic preferences in real scenarios. A larger red area indicates lower model accuracy and poorer discrimination of affected regions. Conversely, the green areas represent predicted labels obtained during the model’s testing phase. A larger green area suggests that the model has mistakenly segmented healthy skin, which could mislead doctors into treating non-affected areas, especially with destructive procedures like lasers or cryotherapy. The yellow areas indicate the overlap between predicted and ground truth labels; a larger yellow area signifies more accurate identification of lesion regions by the model. In summary, from the perspective of clinical diagnosis and treatment, smaller green and red areas and a larger yellow area indicate better model performance, enabling more effective segmentation of skin lesions to assist in diagnosis and treatment decisions.
@@
The qualitative analysis results of the 7 models on the ISIC-2016 dataset are shown in Figure 4(a). From the first two rows, it is evident that ScaleFusionNet exhibits a broader yellow region compared to the D-LKA model, indicating that ScaleFusionNet is better at identifying affected areas. Additionally, ScaleFusionNet’s predictions display fewer green and red regions, which reduces the likelihood of misdiagnosis and missed diagnoses. This difference is even more pronounced in the latter two rows. Although TransFuse has the largest yellow region, it also shows the largest green area, indicating misdiagnosis of healthy regions. While it encompasses lesion areas, such misdiagnosis can have serious implications in clinical diagnosis. In comparison to D-LKA, ScaleFusionNet has a similarly sized yellow region but with a smaller misdiagnosis area, aligning better with clinical diagnostic needs.
@@
These results highlight ScaleFusionNet’s superior ability to accurately segment skin lesions while minimizing errors, making it a highly effective tool for clinical applications. Its combination of high accuracy, and strong generalization ability positions it as a leading solution for skin lesion segmentation tasks. The model’s ability to preserve fine-grained details and refine lesion boundaries further underscores its potential for improving melanoma diagnosis and treatment in real-world clinical settings.
@@ 6 CONCLUSION
This study presents a medical image segmentation model called ScaleFusionNet, which incorporates Cross-Attention Transformer Module (CATM) and AdaptiveFusionBlock. The model utilizes deep learning algorithms to learn and extract complex features from medical images. Specifically, the proposed AdaptiveFusionBlock and CATM significantly enhance performance in skin cancer segmentation tasks. Experiments conducted on publicly available datasets demonstrate that this model achieves competitive results, strongly supporting advancements in skin cancer medical image segmentation. These innovative approaches positively impact diagnostic accuracy, guide treatment decisions, and promote further research in the field. Nonetheless, there are still some issues within this study. While ScaleFusionNet maintains a higher accuracy, its computational complexity is higher compared to some methods. This is primarily due to the integration of multi-scale feature extraction and cross-attention mechanisms, where contributions from different receptive fields may vary. A potential solution could involve optimizing feature allocation in a manner akin to self-selective routing. Additionally, the application of deformable convolutions may exacerbate this issue, making it worthwhile to explore simplifications in deformable operations in future research.
From a clinical application perspective, a trustworthy medical image segmentation method must not only provide high-quality segmentation results but also deliver corresponding uncertainty metrics. ScaleFusionNet represents a significant step forward in skin lesion segmentation, leveraging adaptive multi-scale fusion and cross-attention mechanisms to achieve high accuracy and robustness. However, future work should address computational efficiency and incorporate uncertainty quantification to further enhance its clinical applicability and reliability. By aligning more closely with clinical needs, such as accounting for lesion severity and variability, ScaleFusionNet and its successors can better support dermatologists in diagnosing and treating skin conditions effectively.
@@ Acknowledgements
The authors extend their appreciation to Taif University, Saudi Arabia, for supporting this work through project number (TU-DSPP-2024-229)
@@
Funding Statement: This research was funded by Taif University, Taif, Saudi Arabia project number (TU-DSPP-2024-17).

### v2

@@ Abstract
Melanoma is a malignant tumor originating from skin cell lesions. Accurate and efficient segmentation of skin lesions is essential for quantitative medical analysis but remains a challenge. To address this, we propose ScaleFusionNet, a segmentation model that integrates the Cross-Attention Transformer Module (CATM) and AdaptiveFusionBlock to enhance feature extraction and fusion. The model employs a hybrid architecture encoder that effectively captures both local and global features. We introduce CATM, which utilizes Swin Transformer Blocks and Cross Attention Fusion (CAF) to adaptively refine encoder-decoder feature fusion, reducing semantic gaps and improving segmentation accuracy. Additionally, the AdaptiveFusionBlock is improved by integrating adaptive multi-scale fusion, where Swin Transformer-based attention complements deformable convolution-based multi-scale feature extraction. This enhancement refines the boundaries of the lesion and preserves fine-grained details. ScaleFusionNet achieves Dice scores of 92. 94% and 91. 65% in the ISIC-2016 and ISIC-2018 datasets, respectively, demonstrating its effectiveness in skin lesion analysis. Our code implementation is publicly available at GitHub.
@@ 1 Introduction
The incidence of melanoma has risen significantly in recent decades, driven by increasing environmental pollution and ultraviolet radiation [1]. This trend has drawn significant attention from the global medical community. Early detection is essential for improving treatment outcomes and patient survival. Traditional diagnostic methods, such as clinical observation and tissue biopsy, are limited by subjectivity and invasiveness, which make them unsuitable for large-scale screening. Medical image segmentation offers a noninvasive and high-precision alternative that gives clinicians with more detailed and accurate information. This technology has the potential to greatly enhance early detection and treatment of skin cancer.
@@
With the rapid advancement of deep learning, convolutional neural network (CNN)-based medical image segmentation models have gained prominence, achieving notable success in tasks such as skin lesion segmentation. Among these, the U-Net model, introduced by Ronneberger et al. [2], stands out as a pioneering framework. U-Net has demonstrated remarkable performance in medical image segmentation and has established a distinctive U-shaped architectural paradigm. However, due to the inherent limitations of convolutional operations, CNN-based methods often struggle to model global contextual information effectively. This limitation becomes particularly evident in medical image segmentation tasks with significant inter-sample variability, such as skin lesion segmentation. To address this challenge, researchers have explored various strategies, including the use of large kernel convolutions, dilated convolutions, and other techniques aimed at expanding receptive fields [3, 4, 5, 6, 7]. For example, Hu et al. [8] employed self-attention to enhance readvantageelds, while Tang et al. [9] proposed a model that took adavantage of large convolutional kernels and skip fusion to achieve promising results in tasks such as breast nodule ultrasound image segmentation. Inspired by the ConvNeXt model [10], Han et al. [11] developed a medical image segmentation method based on large kernel convolutions, successfully applying it to tasks such as retinal vessel segmentation. Despite these advancements, simply increasing the size of convolutional kernels may not fully resolve the challenge of modeling global features, as the fundamental constraints of the receptive field remain.
@@
Recently, Transformer [12], which serves as a feature extraction technique with powerful global modeling capabilities, has achieved notable success in both the natural language processing and computer vision domains. Cai et al. [13] unveiled the BiADATU‐Net that combined Transformer and feature adaptation modules, which resulted in promising outcomes in several publicly accessible skin lesion segmentation datasets. Zhang et al. [14] developed DAE-Former, a pure Transformer U-shaped medical image segmentation model, harnessing efficient Transformers steered by dual attention, similar to Swin-Unet [15], and exhibited commendable performance in diverse image segmentation datasets, including ISIC-2018 [16]. However, because Vision Transformers can only output single-scale feature representations, they lack the ability to capture multi-scale information in two-dimensional images [17, 18]. As a result, Transformer-based medical image segmentation models may struggle to seamlessly integrate multi-scale information, leading to insufficient attention to lesion regions and incomplete decoding of feature details.
@@ 5.1 Results on ISIC-2016
To better compare with current mainstream models, 11 prominent models were selected for comparison with the proposed ScaleFusionNet. Table 1 shows that ScaleFusionNet achieved a DSC score of 92.94% and an IOU score of 87.35% on the ISIC-2016 dataset, which are the average results of five repeated experiments, demonstrating outstanding performance. Compared to the Swin-Unet model, ScaleFusionNet improved by 2.82% in the DSC metric and 4.14% in the IOU metric. When compared to MISSFormer, ScaleFusionNet exhibited enhancements of 2.48% and 3.43% in the DSC and IOU metrics, respectively. Against the D-LKA model, ScaleFusionNet still showed improvements of 0.11% and 0.20% in the DSC and IOU metrics, respectively. This indicates that ScaleFusionNet more accurate than D-LKA to detect the refined boundary of skin lesions. Compared to methods like U-Net, although ScaleFusionNet employs a more complex architecture to address potential issues in U-shaped medical image segmentation models, we find this approach justified given the 5.13% performance improvement and the reduced computational load. In clinical applications, a faster and lighter model can support a broader range of compatible use cases, which is crucial for hospitals or organizations with limited computational resources. Concurrently, to enhance the assessment of model performance, we select 7 high-performing models for qualitative scrutiny of experimental outcomes, elucidating the differences among them. The red areas in the illustrations represent ground-truth labels meticulously annotated by experts, reflecting the diagnostic preferences of clinical doctors in real scenarios. A larger red area indicates lower model accuracy and poorer discrimination of affected regions. In contrast, the green areas represent the predicted labels obtained during the model testing phase. A larger green area suggests that the model has mistakenly segmented healthy skin, which could mislead doctors into treating non-affected areas, especially with destructive procedures like lasers or cryotherapy. The yellow areas indicate the overlap between the predicted and ground truth labels; a larger yellow area signifies a more accurate identification of the region of the lesion by the model. In summary, from the perspective of clinical diagnosis and treatment, smaller green and red areas and a larger yellow area indicate better model performance, enabling more effective segmentation of skin lesions to help in diagnosis and treatment decisions.
@@
The results of the qualitative analysis of the 7 models in the ISIC-2016 dataset are shown in Figure 5(a). From the first two rows, it is evident that ScaleFusionNet exhibits a broader yellow region compared to the D-LKA model, indicating that ScaleFusionNet is better at identifying affected areas. Furthermore, ScaleFusionNet’s predictions show fewer green and red regions, which reduces the likelihood of misdiagnosis and missed diagnoses. This difference is even more pronounced in the latter two rows. Although TransFuse has the largest yellow region, it also shows the largest green area, indicating a misdiagnosis of healthy regions. Although it covers areas of injury, this misdiagnosis can have serious implications in the clinical diagnosis. In comparison to D-LKA, ScaleFusionNet has a similarly sized yellow region but with a smaller misdiagnosis area, aligning better with clinical diagnostic needs.
@@
These results highlight ScaleFusionNet’s superior ability to accurately segment skin lesions while minimizing errors, making it a highly effective tool for clinical applications. Its combination of high accuracy, and strong generalization ability positions it as a leading solution for skin lesion segmentation tasks. The model’s ability to preserve fine-grained details and refine the limits of the injury further underscores its potential to improve melanoma diagnosis and treatment in real-world clinical settings.
@@ 5.2 Results on ISIC-2018
Figure  4 visually demonstrates the results of various segmentation models in the ISIC 2016 and 2018 datasets, focusing on both DSC and IOU metrics. As observed, ScaleFusionNet consistently outperforms other models, particularly in the ISIC 2016 dataset, where it achieves the highest DSC score. The blue line for ISIC 2016 DSC shows a steady increase, with ScaleFusionNet clearly marked as the top performer. Similarly, for ISIC 2018 DSC, ScaleFusionNet, indicated by the green line, again achieves a leading position, closely followed by models such as D-LKA and TransFuse. The dashed red and yellow lines represent the IOU scores, where ScaleFusionNet excels in the 2016 dataset, further demonstrating its dominance in multiple evaluation metrics. Overall, the graph illustrates the effectiveness of ScaleFusionNet in both DSC and IOU metrics for both datasets, demonstrating it to be a robust model for skin lesion segmentation tasks.
Figure 4: Performance comparison of segmentation models on ISIC-2016 and ISIC-2018 datasets. The figure shows DSC (circles) and IOU (squares) scores for various models, with ScaleFusionNet outperforming other models in both metrics across both datasets.
@@ 6 CONCLUSION
This study presents a medical image segmentation model called ScaleFusionNet, which incorporates the Cross-Attention Transformer Module (CATM) and AdaptiveFusionBlock. The model utilizes deep learning algorithms to learn and extract complex features from medical images. Specifically, the proposed AdaptiveFusionBlock and CATM significantly enhance performance in skin cancer segmentation tasks. Experiments conducted on publicly available datasets demonstrate that this model achieves competitive results, strongly supporting advances in skin cancer medical image segmentation. These innovative approaches positively impact diagnostic accuracy, guide treatment decisions, and promote further research in the field. However, there are still some issues within this study. While ScaleFusionNet maintains a higher accuracy, its computational complexity is higher compared to some methods. This is primarily due to the integration of multi-scale feature extraction and cross-attention mechanisms, where contributions from different receptive fields may vary. A potential solution could involve optimizing feature allocation in a manner similar to self-selective routing. Additionally, the application of deformable convolutions may exacerbate this issue, making it worthwhile to explore simplifications in deformable operations in future research.
From a clinical application perspective, a reliable medical image segmentation method must not only provide high-quality segmentation results but also deliver corresponding uncertainty metrics. ScaleFusionNet represents a significant step forward in skin lesion segmentation, leveraging adaptive multi-scale fusion and cross-attention mechanisms to achieve high accuracy and robustness. However, future work should address computational efficiency and incorporate uncertainty quantification to further enhance its clinical applicability and reliability. By aligning more closely with clinical needs, such as accounting for the severity and variability of lesions, ScaleFusionNet and its successors can better support dermatologists in diagnosing and treating skin conditions effectively.
@@ Acknowledgements
This work was supported in part by the UK EPSRC under Grant EP/P009727/2, and the Leverhulme Trust under Grant RF-2019-492. The authors also express their appreciation to Taif University, Saudi Arabia, for supporting this work through project number (TU-DSPP-2024-229).
@@
Availability of Data and Materials: The datasets analyzed in the current study are available in the 2016 and 2018 repositories at the following link: https://challenge.isic-archive.com/data/.